### v1

@@ Abstract
The deployment of large language models (LLMs) in healthcare has demonstrated substantial potential for enhancing clinical decision-making, administrative efficiency, and patient outcomes. However, the underrepresentation of diverse groups in the development and application of these models can perpetuate biases, leading to inequitable healthcare delivery. This paper presents a comprehensive scientometric analysis of LLM research for healthcare, including data from January 1, 2021, to June 16, 2024. By analyzing metadata from PubMed and Dimensions, including author affiliations, countries, and funding sources, we assess the diversity of contributors to LLM research. Our findings highlight significant gender and geographic disparities, with a predominance of male authors and contributions primarily from high-income countries (HICs). We introduce a novel journal diversity index based on Gini impurity to measure the inclusiveness of scientific publications. Our results underscore the necessity for greater representation in order to ensure the equitable application of LLMs in healthcare. We propose actionable strategies to enhance diversity and inclusivity in artificial intelligence research, with the ultimate goal of fostering a more inclusive and equitable future in healthcare innovation.
@@ 1 Introduction
The advent of large language models (LLMs) represents a significant breakthrough in artificial intelligence (AI), revolutionizing a wide range of applications across numerous fields[1, 2, 3, 4, 5]. In the context of healthcare, LLMs have demonstrated considerable potential in enhancing clinical decision-making, streamlining administrative processes, and improving patient outcomes through natural language processing tasks such as medical record analysis, automated diagnosis, and personalized treatment recommendations [6]. These advanced models, trained on vast corpora of text data, possess the ability to understand, interpret, and generate human language with remarkable accuracy, thus offering promising solutions to some of the most pressing challenges in modern healthcare [7, 8].
However, the deployment of LLMs in healthcare is not without its challenges, particularly concerning the issue of representation within the machine learning and healthcare communities [9]. Underrepresentation of diverse groups in the development and application of LLMs can lead to the perpetuation or amplification of biases, stereotypes, and negative perceptions of minorities in society [10, 11, 12, 13]. Such biases can adversely affect the quality and equity of healthcare delivery, potentially exacerbating existing disparities [14]. For instance, if LLMs are predominantly developed and trained by researchers from high-income countries (HICs), the resulting models may not adequately address or understand the healthcare needs of populations in low- and middle-income countries (LMICs). This lack of representation can lead to biased algorithms that fail to generalize across different demographic groups, thus limiting the global applicability and fairness of AI-driven healthcare solutions.
@@
In this paper, we present a scientometric review of LLM research in healthcare, utilizing data extracted from PubMed and Dimensions API covering the period from January 1, 2021 to June 16, 2024. Our study involved extracting metadata such as authors’ affiliations, countries, names, and funding information, and estimating the diversity of journals and research in the field of LLMs applied to healthcare. Through this analysis, we aim to shed light on the diversity of contributors to LLM research in healthcare and underscore the need for broader representation to ensure the equitable and effective application of LLMs worldwide. By evidencing the current state of representation and proposing actionable solutions, our study seeks to contribute to the ongoing discourse on diversity in AI and healthcare, ultimately advocating for a more inclusive and equitable future in health technology innovation.
@@ 2.1 Cohort Selection
The cohort for this study was defined through a systematic search of PubMed [19], a premier biomedical literature database, to identify research articles focused on the application of LLMs in healthcare. PubMed was chosen due to its comprehensive coverage of biomedical and life sciences literature, ensuring an extensive collection of relevant studies.
To construct the search query, we used specific terms to capture the essence of LLMs and their application in healthcare. The search terms included:
@@
These terms were designed to be inclusive enough to cover various aspects and terminologies related to LLMs in healthcare. The search was restricted to articles published between January 1, 2021, and June 16, 2024, to capture recent advancements and trends in this rapidly evolving field. The retrieved articles formed the initial dataset for our scientometric analysis. The details of each article, including titles, abstracts, and publication dates, were extracted to facilitate further examination and classification. This initial extraction served as the foundation for subsequent steps in our analysis, where we aimed to gather comprehensive metadata about the authors and their affiliations to assess the diversity of research contributions in this domain.
@@ 2.2 Metadata Extraction and Preprocessing
Following the initial identification of relevant articles through PubMed, the next step involved the extraction of detailed metadata from the Dimensions database [20]. The PubMed ID (PMID) of each article was used to retrieve comprehensive information about the publications, including details on the authors, their affiliations, and funding sources. This extraction process utilized the Dimensions API, which allowed for efficient retrieval of metadata by querying the database with the specific PMIDs of the selected articles.
The data extracted from Dimensions included key information such as the name of the author, their respective affiliations, the position they held in the authoring process, and the countries in which they were based. Additionally, details about the funding organizations and the countries from which they hailed was included. The dataset enabled us to perform a thorough analysis of the geographic and demographic distribution of research contributions. To facilitate this analysis, we performed several preprocessing steps:
  • Gender Identification: The Genderize.io API [21] was employed to infer the gender of the authors based on their first names. This API uses a large database of names to provide probabilistic estimates of gender, allowing us to classify authors as male or female using the maximum probability.
@@
The preprocessing of the extracted data entailed the cleansing and organization of the data to ensure its accuracy and consistency. This included the handling of missing values, the standardizing country names, and the conversion of country codes to their respective ISO alpha-3 formats, which was achieved through the use of the ‘pycountry‘ library. Additionally, the authors’ names and affiliations were parsed to extract relevant information, and the data was structured to facilitate subsequent analysis.
@@ 2.3 Data Analysis
The data analysis conducted in this study was aimed at evaluating the diversity and representation in LLM research in healthcare, focusing on gender distributions, geographic distributions by income and continent, funder analysis, and journal diversity.
Gender Distributions: A gender estimation was conducted using the Genderize.io API, which classified genders according the methodology described in the previous section. Using the gender proportion, the number of female authors across the dataset and the position in the paper was used to calculate the overall proportions.
@@
Continent and Income Group Distributions: From the Authors’ information extracted in the previous section grouped by continents (Africa, Asia, Europe, North America, Oceania, South America) and classified into HIC and LMIC, The distribution of authors by continent and income group was analyzed to highlight geographic disparities in LLM research contributions.
Funding Source Analysis: The funding source analysis involved the identification of the countries from which the funding organizations resided, and categorizing them into HICs and LMICs. The proportions of publications funded by organizations from each income group were calculated in order to identify potential biases in research funding.
@@
Distribution of Publications per Country: The distribution of publications per country was analyzed by aggregating the number of publications attributed to authors from each country. A choropleth map was employed to illustrate the geographic distribution of research contributions.
@@
Journal Diversity and Index Testing: Journal diversity was assessed using a novel index based on Gini impurity, measuring the conditional probabilities of an author being of a certain gender, from a specific country, and from a particular income group, given that they are an author in a specific journal. The diversity index $D_{j}$ for a journal $j$ is calculated as:
$D_{j}=1-\sum_{i=1}^{n}\left[P(G=i\mid J=j)\cdot P(C=i\mid J=j)\cdot P(I=i\mid J=j)\right]^{2}$
@@
The index was validated on journals with more than 100 authors to ensure robustness. This metric is a key contribution of our analysis, providing a quantitative measure of diversity in scientific publishing.
Overall, the data analysis yielded a comprehensive assessment of diversity and representation in LLM research in healthcare, identifying areas for improvement and proposing measures to enhance inclusivity in this critical field.
@@ 3.2 Gender Distributions
The results showed disparities in gender representation across different author positions (overall, first authors, and last authors), and these distributions were further categorized by HICs and LMICs. Tables 1, 2, and 3 summarize the different gender distributions.
Table 1: Overall Gender Distribution of authors in the Dataset.
Author Position Male (%) Female (%)
Overall 76.29 23.71
First Authors 79.61 20.39
Last Authors 83.40 16.60
Table 2: Gender Distribution of atuhors in HIC papers
Author Position Male (%) Female (%)
Overall 76.55 23.45
First Authors 79.31 20.69
Last Authors 84.38 15.62
Table 3: Gender Distribution of authors in LMIC papers
Author Position Male (%) Female (%)
Overall 74.87 25.13
First Authors 81.48 18.52
Last Authors 76.67 23.33
@@ 3.3 Distribution of Papers per Continent for LMICs and HICs
The distribution of papers was analyzed by continent and income group (HICs and LMICs). The results are summarized in the Table 4. Additionally, a world map illustrating the distribution of research contributions across continents provides a geographic perspective on paper diversity. (Figure 1)
Table 4: Distribution of Papers by Continent and Country Income Group
Continent Income Group N Proportion (%)
Africa LMIC 26 0.29
Asia HIC 703 7.77
Asia LMIC 1066 11.79
Europe HIC 1808 19.99
Europe LMIC 21 0.23
North America HIC 5197 57.46
Oceania HIC 176 1.95
South America HIC 10 0.11
South America LMIC 38 0.42
@@
Figure 2: Distribution of Papers Published by Continent and Income Group
@@ 3.4 Funding Source Analysis
The analysis of funding sources was categorized by continent and income group. The results indicate potential biases in the distribution of funding across different regions. Table 5 summarizes these findings.
Table 5: Distribution of Funding Countries by Continent and Income Group
Continent Income Group N Proportion (%)
Africa LMIC 1 0.21
Asia HIC 39 8.25
Asia LMIC 52 10.99
Europe HIC 88 18.61
North America HIC 286 60.47
Oceania HIC 4 0.85
South America HIC 1 0.21
South America LMIC 2 0.43
Figure 3: Proportion of Funding Source by Continent and Income Group

### v2

@@ Abstract
The deployment of large language models (LLMs) in healthcare has demonstrated substantial potential for enhancing clinical decision-making, administrative efficiency, and patient outcomes. However, the underrepresentation of diverse groups in the development and application of these models can perpetuate biases, leading to inequitable healthcare delivery. This paper presents a comprehensive scientometric analysis of LLM research for healthcare, including data from January 1, 2021, to July 1, 2024. By analyzing metadata from PubMed and Dimensions, including author affiliations, countries, and funding sources, we assess the diversity of contributors to LLM research. Our findings highlight significant gender and geographic disparities, with a predominance of male authors and contributions primarily from high-income countries (HICs). We introduce a novel journal diversity index based on Gini diversity to measure the inclusiveness of scientific publications. Our results underscore the necessity for greater representation in order to ensure the equitable application of LLMs in healthcare. We propose actionable strategies to enhance diversity and inclusivity in artificial intelligence research, with the ultimate goal of fostering a more inclusive and equitable future in healthcare innovation.
@@ 1 Introduction
The advent of large language models (LLMs) represents a significant breakthrough in artificial intelligence (AI), revolutionizing a wide range of applications across numerous fields[1, 2, 3, 4, 5]. In the context of healthcare, LLMs have demonstrated considerable potential in enhancing clinical decision-making, streamlining administrative processes, and improving patient outcomes through natural language processing tasks such as medical record analysis, automated diagnosis, and personalized treatment recommendations [6]. These advanced models, trained on vast corpora of text data, possess the ability to understand, interpret, and generate human language, thus offering promising solutions to some of the most pressing challenges in modern healthcare [7, 8].
However, the deployment of LLMs in healthcare is not without its challenges, particularly the need of data and computational resources [9], concerning the issue of representation within the machine learning and healthcare communities [10]. Underrepresentation of diverse groups in the development and application of LLMs can lead to the perpetuation or amplification of biases, stereotypes, and negative perceptions of minorities in society [11, 12, 13, 14]. Such biases can adversely affect the quality and equity of healthcare delivery, potentially exacerbating existing disparities [15, 16]. For instance, if LLMs are predominantly developed and trained by researchers from high-income countries (HICs), the resulting models may not adequately address or understand the healthcare needs of populations in low- and middle-income countries (LMICs). This lack of representation can lead to biased algorithms that fail to generalize across different demographic groups, thus limiting the global applicability and fairness of AI-driven healthcare solutions. Similarly, researchers from HICs are likely to have distinct priorities and blind spots when determining research and developing directions, which may not serve the most pressing needs of diverse populations.
@@
In this paper, we present a scientometric review of LLM research in healthcare, using data extracted from PubMed and Dimensions API covering the period from January 1, 2021 to July 1, 2024. Through this analysis, we aim to shed light on the diversity of contributors to LLM research in healthcare and underscore the need for broader representation to ensure the equitable and effective application of LLMs worldwide. By evidencing the current state of representation and proposing actionable solutions, our study seeks to contribute to the ongoing discourse on diversity in AI and healthcare, ultimately advocating for a more inclusive and equitable future in health technology innovation.
@@ 2.1 Cohort Selection
The cohort for this study was defined through a systematic search of PubMed [21] to identify research articles focused on the application of LLMs in healthcare. PubMed was chosen due to its comprehensive coverage of biomedical and life sciences literature, ensuring an extensive collection of relevant studies.
To construct the search query, we used specific terms to capture studies of LLMs and their application in healthcare. The search terms included:
@@
These terms were designed to be inclusive enough to cover various aspects and terminologies related to LLMs in healthcare. The search was restricted to articles published between January 1, 2021, and July 1, 2024, to capture recent advancements and trends in this rapidly evolving field. The retrieved articles formed the initial dataset for our scientometric analysis. The details of each article, including titles, abstracts, and publication dates, were extracted to facilitate further examination and classification. This initial extraction served as the foundation for subsequent steps in our analysis, where we aimed to gather comprehensive metadata about the authors and their affiliations to assess the diversity of research contributions in this domain.
@@ 2.2 Metadata Extraction and Preprocessing
Following the initial identification of relevant articles through PubMed, the next step involved the extraction of detailed metadata from the Dimensions database [22]. The PubMed ID (PMID) of each article was used to retrieve information about the publications, including details on the authors, their affiliations, and funding sources. This extraction process utilized the Dimensions API, which allowed for efficient retrieval of metadata by querying the database with the specific PMIDs of the selected articles.
Details about the funding organizations and the countries of the funder organizations were included. To facilitate the analysis, we performed several preprocessing steps:
  • Gender Identification: The Genderize.io API [23] was employed to infer the gender of the authors based on information such as their first names, last names, and country. This API uses a large database of names to provide probabilistic estimates of gender, allowing us to classify authors as male or female using the maximum probability.
@@
The preprocessing of the extracted data entailed the cleaning and formatting of the data to ensure its accuracy and consistency. The authors’ names and affiliations were parsed to extract relevant information such as the inference of sex and country from the affiliation country. To standardize the country names, a conversion of country codes to their respective ISO alpha-3 formats was performed. The country formatting conversion was performed through the use of the ‘pycountry‘ library. To facilitate the analysis, we also grouped the countries into continents, having: Africa, Asia, Europe, North America, Oceania, and South America.
@@ 2.3 Data Analysis
In order to archive a global analysis of different variables, we conducted analysis of different variables:
2.3.1 Gender Distributions
The number of male and female authors across the dataset and their position in the author list was used to calculate overall proportions. The overall proportions of male and female authors are calculated as follows:
Let $N_{\text{male}}$ and $N_{\text{female}}$ represent the total counts of male and female authors, respectively, in $D$. The overall proportions $P_{\text{male}}$ and $P_{\text{female}}$ are given by:
$P_{\text{male}}=\frac{N_{\text{male}}}{N_{\text{male}}+N_{\text{female}}},\quad P_{\text{female}}=\frac{N_{\text{female}}}{N_{\text{male}}+N_{\text{female}}}$ (1)
For specific author positions $p$ (e.g., first author or last author), the proportions $P_{\text{male},p}$ and $P_{\text{female},p}$ are similarly calculated:
$P_{\text{male},p}=\frac{N_{\text{male},p}}{N_{\text{male}}+N_{\text{female}}},\quad P_{\text{female},p}=\frac{N_{\text{female},p}}{N_{\text{male}}+N_{\text{female}}}$ (2)
Finally, to compare these proportions with the overall distribution, we use a proportions z-test [25].
2.3.2 Continent and Income Group Distributions per author and funding source
@@
Author and Funding country affiliations were grouped by continents (Africa, Asia, Europe, North America, Oceania, South America) and classified into HIC and LMIC, The distribution of authors and funding sources by continent and income group was analyzed to highlight geographic disparities in LLM research contributions.
The analysis involved grouping papers based on the continent and income group associated with the authors’ affiliations and also with the funding sources. The analysis for authors and income groups was performed independently, however, the calculation method was identical.
The distributions were calculated by dividing the number of papers having authors or funding sources from each continent-income group pair by the total number of papers analyzed. The proportion for each group was calculated as:
$P_{i,j}=\frac{N_{i,j}}{\sum_{i,j}N_{i,j}}$ (3)
Where:
  • $P_{i,j}$ is the proportion of papers having authors or funding sources for continent $i$ and income group $j$.
  • $N_{i,j}$ is the count of papers authors or funding sources for continent $i$ and income group $j$.
  • The summation $\sum_{i,j}N_{i,j}$ represents the total number of papers across all continent-income group pairs.
To validate the robustness of our results, we conducted a sensitivity analysis using a bootstrap sampling method [26]. Let $D$ be the original dataset, where $D\in\mathbb{R}^{n}$ and $n$ represents the total number of observations. The bootstrap method involved the generation of multiple resampled datasets, each denoted as $D_{k}$, where $k$ ranges from 1 to $K$. In our case k was set to be 1000.
For each iteration $k$, the dataset $D_{k}$ is created by randomly sampling with replacement from the original dataset $D$. This process ensures that $D_{k}$ retains the same dimensionality as $D$, but with potentially different distributions due to the resampling. The resampling process can be expressed as:
$D_{k}=\{d_{1}^{k},d_{2}^{k},\dots,d_{n}^{k}\}\quad\text{where}\quad d_{i}^{k}\sim D,\quad i=1,\dots,n$ (4)
Once the resampled dataset $D_{k}$ is generated, the proportion of papers with authors or founding sources for each continent and income group $P_{i,j}^{k}$ is calculated using the equation 3.
After performing $K$ iterations, the results are aggregated to obtain the mean proportion $\hat{P}_{i,j}$ across all bootstrap samples. Additionally, to assess the variability and confidence of the estimates, a 95% confidence interval is computed for each proportion. The mean proportion and confidence intervals are given by:
$\hat{P}_{i,j}=\frac{1}{K}\sum_{k=1}^{K}P_{i,j}^{k}$ (5)
$\text{CI}_{95\%}(P_{i,j})=\left[\text{Percentile}_{2.5\%}(P_{i,j}^{k}),\text{Percentile}_{97.5\%}(P_{i,j}^{k})\right]$ (6)
where $\text{Percentile}_{x\%}(P_{i,j}^{k})$ denotes the $x$-th percentile of the bootstrap distribution for the proportion $P_{i,j}^{k}$.
2.3.3 Distribution of Publications per Country:
@@
The distribution of publications per country was analyzed by aggregating the number of publications attributed to authors from each country. A choropleth map was used to illustrate the geographic distribution of research contributions.
2.3.4 Journal Diversity and Index Testing:
@@
Journal diversity was assessed using an index based on Gini impurity or gini diversity index [27], measuring the conditional probabilities of an author being of a certain gender, from a specific continent, and from a particular income group, given that they are an author in a specific journal. The inclusion of income group and continent was to avoid HIC countries skewing the distributions per continents. Then, the diversity index $D_{j}$ for a given journal $j$ is calculated using the equation:
$D_{j}=1-\sum_{i=1}^{n}\left[P(G=i\mid J=j)\cdot P(C=i\mid J=j)\cdot P(I=i\mid J=j)\right]^{2}$ (7)
@@
The index was calculated on journals where more than 50 authors published LLM papers. This metric is a key contribution of our analysis, providing a quantitative measure of diversity in scientific publishing.
@@ 3.2 Gender Distributions
The results showed disparities in gender representation across different author positions (overall, first authors, and last authors). As can be seen in table 1a, female authors tend to be less represented in the literature representing only 23.84% of the authors compared with 76.16%. The distributions are even more skewed for first and last authors with only 3.06% and 2.71% female authors respectively, compared with 11.90% and 14.04% respectively. As can be seen in table 1b and 1c, these distributions were further categorized by HICs authors and LMICs authors finding similar patterns.
Table 1: Gender Distribution Among Authors by Region
Category Female (Proportion %) Male (Proportion %) P-Value
Table 1a: Distribution Over All Authors
Overall 1963 (23.84%) 6271 (76.16%)
First author, n (%) 252 (3.06%) 980 (11.90%) $<$0.005
Last author, n (%) 223 (2.71%) 1156 (14.04%) $<$0.001
Table 1b: Distribution Over HIC Authors
Overall 1703 (23.55%) 5527 (76.45%)
First author, n (%) 216 (2.99%) 835 (11.55%) $<$0.05
Last author, n (%) 185 (2.56%) 1027 (14.20%) $<$0.001
Table 1c: Distribution Over LMIC Authors
Overall 262 (25.27%) 775 (74.73%)
First author, n (%) 37 (3.57%) 155 (14.95%) 0.0342
Last author, n (%) 38 (3.66%) 137 (13.21%) 0.2357
@@ 3.3 Distribution of Papers per Continent for LMICs and HICs
The analysis of the distribution of research papers across continents and income groups (HICs and LMICs) reveals significant disparities. As shown in Table 2, accounts for the majority of contributions from HICs, with 57.07% of the total papers. Europe follows, contributing 20.17% from HICs. In contrast, contributions from LMICs are markedly lower, with Asia leading at 11.90%, driven largely by contributions from China. Africa, Oceania, and South America contribute minimally to the global research output, especially among LMICs.
The world map in Figure 1 provides a geographic visualization of these contributions, further highlighting the dominance of the United States and China in global research, potentially skewing the overall representation of both HICs and LMICs.
Africa’s contribution remains minimal, despite representing a significant portion of the world’s LMICs, reflecting a broader issue of underrepresentation in global research outputs from this continent.
Table 2: Distribution of Papers by Continent and Income Group
Continent LMIC (Proportion %) HIC (Proportion %)
Africa 31 (0.31%) 0 (0.00%)
Asia 1172 (11.90%) 774 (7.86%)
Europe 32 (0.32%) 1986 (20.17%)
North America 0 (0.00%) 5620 (57.07%)
Oceania 0 (0.00%) 183 (1.86%)
South America 40 (0.41%) 10 (0.10%)
@@
The bootstrap sensitivity analysis, presented in Table 3, further confirms these findings. North America’s dominance in the HIC category is robust across different samples, with a mean proportion of 57.11% (95% CI: 53.95%, 60.10%). Europe’s contribution remains substantial as well, with a mean proportion of 20.19% (95% CI: 17.56%, 22.75%).
On the other hand, the contributions from LMICs, particularly in Africa and South America, are consistently low, even when accounting for sampling variability. The mean proportion for Africa is only 0.33% (95% CI: 0.10%, 0.71%), and for South America, it is 0.41% (95% CI: 0.10%, 0.91%). These results reinforce the observation that the distribution of research output is heavily skewed towards certain high-contributing regions, with minimal representation from others.
Table 3: Bootstrap Sensitivity Analysis of Paper Proportions by Continent and Income Group (%)
Continent LMIC (Mean, CI) HIC (Mean, CI)
Africa 0.33% (0.10%, 0.71%) 0.00% (0.00%, 0.00%)
Asia 11.91% (9.91%, 13.91%) 7.80% (6.10%, 9.46%)
Europe 0.34% (0.10%, 0.71%) 20.19% (17.56%, 22.75%)
North America 0.00% (0.00%, 0.00%) 57.11% (53.95%, 60.10%)
Oceania 0.00% (0.00%, 0.00%) 1.85% (1.11%, 2.75%)
South America 0.41% (0.10%, 0.91%) 0.16% (0.10%, 0.40%)
@@ 3.4 Funding Source Analysis
The raw distribution of funding sources (table 4), shows that North America, particularly within the HIC group, dominates the global research funding landscape, contributing 59.15% of the total funding. Europe follows with 20.42% of the funding sources, also concentrated in HICs. Asia contributes 11.18% from LMICs and 7.90% from HICs, while contributions from other regions such as Africa and South America remain minimal, particularly within the LMIC group.
Table 4: Distribution of Funding Sources by Continent and Income Group
Continent LMIC (Proportion %) HIC (Proportion %)
Africa 1 (0.19%) 0 (0.00%)
Asia 58 (11.18%) 41 (7.90%)
Europe 0 (0.00%) 106 (20.42%)
North America 0 (0.00%) 307 (59.15%)
Oceania 0 (0.00%) 3 (0.58%)
South America 2 (0.39%) 1 (0.19%)
The bootstrap sensitivity analysis in table 5 confirms North America’s dominance in funding is consistently observed, with a mean proportion of 59.73% (95% CI: 46.67%, 72.09%) for HICs. Europe’s contribution remains significant, with a mean proportion of 19.98% (95% CI: 9.30%, 31.38%) for HICs, reflecting a stable pattern across different resampling scenarios.
In contrast, the contributions from LMICs, particularly in Africa and South America, show greater variability but remain low overall. For instance, Africa’s mean proportion of funding from LMICs is 2.00% (95% CI: 1.64%, 3.25%), while South America’s LMIC contribution is 2.12% (95% CI: 1.60%, 3.80%).
Table 5: Bootstrap Sensitivity Analysis of Funding Source Proportions by Continent and Income Group (%)
Continent LMIC (Mean, CI) HIC (Mean, CI)
Africa 2.00% (1.64%, 3.25%) 0.00% (0.00%, 0.00%)
Asia 11.11% (3.70%, 20.41%) 7.92% (1.87%, 15.83%)
Europe 0.00% (0.00%, 0.00%) 19.98% (9.30%, 31.38%)
North America 0.00% (0.00%, 0.00%) 59.73% (46.67%, 72.09%)
Oceania 0.00% (0.00%, 0.00%) 2.18% (1.61%, 4.00%)
South America 2.12% (1.60%, 3.80%) 2.04% (1.56%, 3.74%)